OpenStudy (anonymous):

Trying to answer a compound interest question, Don't need the answer just a little clarification in the steps. 6000=4000(1+r)^5 ?

OpenStudy (campbell_st):

so are you having to find r...?

OpenStudy (wolf1728):

Seems as if the person who asked the question didn't stick around for the answer.

OpenStudy (anonymous):

Im sorry, yes I am, but I wasn't sure if my steps were correct. Let me post.

OpenStudy (wolf1728):

Well I have the formula to solve for rate log(1 + rate) = {log(total) -log(Principal)} ÷ Years Okay - welcome back

OpenStudy (campbell_st):

ok

OpenStudy (anonymous):

the question was at what interest rate (compounded annually) will a sum of 4000$ grow to 6000$ in 5 years.

OpenStudy (campbell_st):

so do you have any working...?

OpenStudy (anonymous):

so I have to solve for r but, I don't know what im doing, so I get 6000=4000(1+r)^5 /4000 /4000 1.5= (1+r)^5 1.5^5/2= 1+r ?

OpenStudy (campbell_st):

well you are correct at 1.5 = (1 + 5)^5 so take the 5th root of both sides of the equation, if you don't know logs then you have \[\sqrt[5]{1.5} = 1 + r\] now solve for r, which will be a decimal, then multiply by 100 to get a percentage

OpenStudy (campbell_st):

oops should be 1.5 = (1 +r)^5

OpenStudy (anonymous):

ohhh.... ok. so if otherwise if I were to use logs which step would I do that. ?

OpenStudy (wolf1728):

4,000 to 6,000 in 5 years log(1 + rate) = {log(total) -log(Principal)} ÷ Years log (1+rate) = [log (6,000) -log(4,000)] / 5 log (1+rate) = [3.7781512504 -3.6020599913] / 5 log (1+rate) = 0.1760912591 / 5 log(1+rate) = 0.0352182518 10^.0352182518 = 1.0844717712 = 1 + rate rate = .0844717712 or 8.44717712%

OpenStudy (anonymous):

i'm sorr@wolf thank you for the help as well, wish I cold give everyone medals.

OpenStudy (campbell_st):

my suggestion, its easier without logs..

OpenStudy (anonymous):

thank you.

OpenStudy (wolf1728):

Yes it is easier without logs :-)